# A Classification Theorem for Steady Euler Flows

### v1

@@ 1.4.3 Moving Plane Arguments
We may also assume that $F(s)\approx\sqrt{s}$ as $s\rightarrow 0.$ As mentioned above, it is not difficult to show that $\psi$ cannot change sign in $\Omega_{\Gamma}$. By scaling, we can assume that $0\leq\psi\leq 1$ Now, the overdetermined problem above seems similar to those considered by Serrin in his landmark paper [39]. Unfortunately, a key difference here is that we do not have the Lipschitz smoothness of the non-linearity $F,$ which is very important for the moving plane method used in [39]. Indeed, as we have already said, near $s=0,$ $F(s)\approx\sqrt{s}$; moreover, $F$ need not be smooth as $s\rightarrow 1$ (the maximum value of $\psi$) either. To deal with the singularity as $s\rightarrow 0,$ we use the fact that we have a lower bound on the growth of $\psi$ in the direction normal to the boundary, since the vorticity is non-zero (and constant) on the boundary. Then we invoke a refinement of the Hopf lemma for elliptic problems on small domains under some growth conditions on the coefficients. This effectively deals with the singularity as $s\rightarrow 0$. Remarkably, it turns out that the potential singularity as $s\rightarrow 1$ always has the right sign for the maximum principle (see the proof of Lemma 4.11).
1.5 Organization of the Paper
In Section 2, we establish some general results about analytic functions satisfying $\{f,g\}=0$ and recall a few results about the zero sets of analytic functions. The main consequence of this is that, if an analytic Euler solution $\psi$ does not satisfy an equation of the type (1.4), we are ensured the existence of a level set of $\psi$ on which the gradient vanishes. In Section 3, we prove a few results about these sets, using more fundamentally that $\psi$ is an Euler solution. In particular, we show the analyticity of the non-isolated portions of the critical points with quadratic vanishing. This sets us up with an overdetermined elliptic problem in a subdomain of the original domain. At the end of that section, we then establish the rigidity of solutions with curves of critical points with higher than quadratic vanishing. In Section 4, we establish the rigidity of solutions with a suitably chosen curve of critical points with quadratic vanishing first by deducing certain regularity properties of the function $F$ (defined in a suitable subdomain) and then by applying the moving plane method. In Section 4.3 we give a counterexample to Theorem 1.1 in an analytic non-simply connected bounded domain, hence showing the optimality of the statement.
2 Facts about $\{f,g\}=0$
@@
2.1 Key Facts
We start by giving a notation for the degree of flatness of an analytic function at a point.
@@ 2.1 Key Facts
For any non-constant analytic $f:\Omega\rightarrow\mathbb{R},$ and $x_{*}\in\Omega,$ we define $d_{f}(x_{*})$ to be the first $k\geq 1$ for which $D^{k}(f)(x_{*})\not=0.$
We then introduce notation for the set of critical points of $f$ and those with even vanishing:
@@
Assume that $\{f,g\}=0$ in some open connected set $\Omega\subset\mathbb{R}^{2}.$ Suppose that $\Gamma$ is an analytic curve connecting $x,y\in\Omega$ and that $\Gamma\cap\mathcal{C}_{e}(f)=\emptyset.$ Then, if $f(x)=f(y),$ it must be that $g(x)=g(y).$
To prove this proposition, let us first establish a generalization of the well-known straightening lemma.
@@
The case $d=1$, which is that $\nabla f(x_{*})\not=0,$ is the standard straightening lemma which is valid only under the assumptions that $f$ and $g$ are $C^{1}$ in a neighborhood of $x_{*}.$
@@
Strictly speaking, we do not need the generalization to $d>1$ for the results proven later; however, it is conceptually important in order to understand the difference between “extendible” and “unextendible” domains later. The result may also be of independent interest. It should be noted that, in this generalization, analyticity plays a key role with the statement not holding true in the smooth class.
Proof.
Without loss of generality, by applying suitable translations and a rotation, set $x_{*}=0$, $f(x_{*})=0,$ and $\partial_{1}^{d}f(0)>0.$ Assume that $d$ is odd. The point is that we have that $\partial_{1}f(t,0)>0$ for $t\in[-\epsilon,0)\cup(0,\epsilon]$ for some $\epsilon>0$. In particular, the function $f(t,0)$ is real-analytic and invertible in $(-\epsilon,\epsilon)$. It follows that we can express its inverse using a Puiseux series for $t$ sufficiently small. In particular, there exists (a unique) $F$ that is analytic in the variable $s^{1/d}$ for which
$g(t,0)=F(f(t,0)),$
for every $t\in(-\frac{\epsilon}{2},\frac{\epsilon}{2}).$ Let $\kappa$ be the radius of convergence of the Puiseux series of $F$ at $0$ We now take $\delta<\frac{\epsilon}{2}$ be so small that $f(B_{\delta}(0))\subset(-\kappa,\kappa).$ We will show that
$g(x)=F(f(x)),$
for every $x\in B_{\delta}(0).$ Fix any $x\in B_{\delta}(0).$ First, take $x_{1}\in(-\delta,\delta)\setminus\{0\}$. By assumption, $(x_{1},0)$ is not a critical point of $f$. It follows that there is a real analytic $G$ so that $g(x)=G(f(x)),$ for every $x$ in a neighborhood of $(x_{1},0).$ In particular, there is an $\epsilon(x_{1})>0$ such that for all $t\in(x_{1}-\epsilon(x_{1}),x_{1}+\epsilon(x_{1})),$ we have that
@@
At this point, we have shown that there is a ball around each point of the segment $\{(t,0):t\in(-\epsilon,\epsilon)\setminus\{0\}\}$ where $g=F(f).$ Now we will extend this to the whole ball $B_{\delta}(0).$ Fix any $x\in B_{\delta}(0)$ with $x_{1}\neq 0$. Without loss of generality, assume $x_{2}>0.$ Now consider the function $g(x_{1},t),$ where $t\in[0,x_{2}].$ This is an analytic function of $t$, since $g$ is an analytic. Now, $F(f(x_{1},t))$ is analytic and equal to $g$ in a neighborhood of $t=0$. Moreover, $F(f(x_{1},t))$ is also analytic except at the possibly discrete set of points where $f(x_{1},t)=0$, where $F(f(x_{1},t))$ has a convergent Puiseux series (note that $f(x_{1},0)\not=0$). It follows that $F(f(x_{1},t))$ is actually analytic on $[0,x_{2}]$ and that $g(x_{1},t)=F(f(x_{1},t))$ for every $t\in[0,x_{2}].$ In particular, $g(x_{1},x_{2})=F(f(x_{1},x_{2})).$
@@
For any point $x\in\Gamma$, by applying Lemma 2.3 to $f$ and $g$, we can find a neighborhood $U_{x}$ and a Puiseux function $F_{x}$ such that $g=F(f)$ in $U_{x}$. As the collection of the open set $U_{x}$ forms an open cover for the compact set $\Gamma$, we can choose $U_{x^{1}},\cdots,U_{x^{m}}$, such that $U_{x^{i}}\cap U_{x^{i+1}}\neq\emptyset$ and $\Gamma\subset\cup_{i=1}^{m}U_{x^{i}}$. In the case where $f$ is constant in $\Gamma$, as the pair $(g,f)$ satisfies a function relationship in each $U_{x^{i}}$, we have $g$ is constant in $U_{x^{i}}\cap\Gamma$. Notice that $\{U_{x^{i}}\}$ is a finite open cover for $\Gamma$, we have $g$ is constant in $\Gamma$.
In the case where $f$ is not constant in $\Gamma$, we claim $F_{x^{1}}$ can be extended to be a Puiseux function $F$ in the range of $f([\Gamma])$ with $g=F(f)$ in $\Gamma$. We first prove that $F_{x^{1}}$ can be extended to a Puiseux function $F$ in the range $f(U_{x^{1}}\cap\Gamma)\cup f(U_{x^{2}}\cap\Gamma)$, with $g=F(f)$ in $\Gamma\cap(U_{x^{1}}\cup U_{x^{2}})$. Here $F_{x^{1}}$ and $F_{x^{2}}$ are two Puiseux functions respectively in the range of $f(U_{x^{1}}\cap\Gamma)$ and $f(U_{x^{2}}\cap\Gamma)$, and the values of $F_{x^{1}}$ and $F_{x^{2}}$ agree in $f(\Gamma\cap U_{x^{1}}\cap U_{x^{2}})$. As $f$ is not constant in $\Gamma$, the analytic assumption of $f$ and $\Gamma$ implies that $f(\Gamma\cap U_{x^{1}}\cap U_{x^{2}})$ contains an open interval in $\mathbb{R}$. Now due to the fact that $f(\Gamma\cap U_{x^{1}})$ and $f(\Gamma\cap U_{x^{2}})$ are connected, the condition that $F_{x^{1}}$ and $F_{x^{2}}$ agree in $f(U_{x^{1}}\cap U_{x^{2}})$ implies that the values of $F_{x^{1}}$ and $F_{x^{2}}$ agree in $f(\Gamma\cap U_{x^{1}})\cap f(\Gamma\cap U_{x^{2}})$. Hence, we can extend $F_{x^{1}}$ with the aid of $F_{x^{2}}$ to a Puiseux function $F$ in $f(U_{x^{1}})\cup f(U_{x^{1}})$ and $g=F(f)$ in $\Gamma\cap(U_{x}^{1}\cup U_{x}^{2})$. We can iterate the above procedure to extend $F_{x^{1}}$ to a Puiseux function $F$ in the range of $f([\Gamma])$ such that $g=F(f)$ in $\Gamma$ and we now finish the proof. ∎
2.2 Zero Sets of Analytic Functions
In this section, we will briefly discuss the structure of zero set of a real analytic function in two variables. We firstly recall a classical result concerning the structure of zero set of real analytic function.
@@ 2.2 Zero Sets of Analytic Functions
  3. 3.
    The nodal set $\{f=0\}$ near $x^{0}$ is a collection of an even number of analytic curves branching at $x^{0}$.
It is an immediate corollary of Proposition 2.6 that for any given point $x$, which is not isolated in the nodal set for a non-constant analytic function $f$, there are at least two analytic curves $\Gamma_{1}$, $\Gamma_{2}$ branching at $x$ and consisting of nodal points of $f$. In most cases, we cannot glue $\Gamma_{1}$ and $\Gamma_{2}$ together to form a single analytic curve. For example, in the case $f(x,y)=x^{2}-y^{3}$, the nodal set has a cusp at the origin. However, in the case where $\Gamma_{1}$ consists of critical points of the same degree (in particular, if the degree doesn’t jump up at the common point), $\Gamma_{1}\cup\Gamma_{2}$ is an analytic curve.
@@
Let $x^{0}$ be a fixed point and we assume there is an analytic curve $\Gamma_{1}:[0,1]\rightarrow\mathbb{R}^{2}$ with $\Gamma_{1}(0)=x^{0}$, and $d_{f}(x)=m$ for all $x\in\Gamma_{1}([0,1])$. Then the level set $\{x|f(x)=f(x^{0})\}$ near $x^{0}$ is an analytic curve $\Gamma$. Moreover, for any point $x\in\Gamma$, $d_{f}(x)=m$.
@@
Without loss of generality, by applying suitable rotation, we assume $x^{0}=0$, $f(0)=0$ and $\partial_{1}^{m}f(0)>0$. By applying the analytic inverse function theorem to $\partial_{1}^{m-1}f$ at 00, we have the set $\{\partial_{1}^{m-1}f=0\}$ near 00 is an analytic curve $\Gamma$ passing through 00. The existence of the analytic curve $\Gamma_{1}$ consisting of critical points of degree $m$ that starts at 00 ensures that for infinitely many points $x^{i}$ accumulate at 00 in $\Gamma$, we have
$D^{k}f(x^{i})=0,\text{for $1\leq k\leq m-1$ and $i$. }$ (2.1)
As $\Gamma$ and $f$ are analytic, we have
$D^{k}f(x)=0,\text{for $x\in\Gamma$, and $1\leq k\leq m-1$.}$ (2.2)
Then due to the fundamental theorem of calculus, we have $f=0$ in $\Gamma$.
Now it suffices to find an open set $\mathcal{O}$ near the origin such that $\Gamma\cap\mathcal{O}$ consists of critical point of degree $m$ and $\{f=0\}\cap\mathcal{O}=\Gamma\cap\mathcal{O}$. We notice the condition $\partial_{1}^{m}f(0)>0$ ensures that $(1,0)$ is not a tangent vector of $\Gamma$ at 00. Now as $f$ and $\Gamma$ are analytic, we can choose $\mathcal{O}:=(-\epsilon,\epsilon)^{2}$ to be a small cube near 00 with the requirement that all the points $x$ in $\mathcal{O}$ satisfy $\partial_{1}^{m}f>0$ and in addition, for any point $x\in\mathcal{O}\setminus\Gamma$, there is a horizontal line $l_{x}$ connecting $x$ and a point $x_{\Gamma}$ in $\Gamma$. Now as $\partial_{1}f(x_{\Gamma})=\cdots=\partial_{1}^{m-1}f(x_{\Gamma})=0$ and $\partial_{1}^{m}f>0$ in $l_{x}$, by applying mean value theorem to $f$ in $l_{x}$, we have $f(x)\neq 0$. ∎
3 Structure of the set of critical points

### v2

@@ 1.4.3. Moving Plane Arguments
We may also assume that $F(s)\approx\sqrt{s}$ as $s\rightarrow 0.$ As mentioned above, it is not difficult to show that $\psi$ cannot change sign in $\Omega_{\Gamma}$. By scaling, we can assume that $0\leq\psi\leq 1.$ Now, the overdetermined problem above seems similar to those considered by Serrin in his landmark paper [46]. Unfortunately, a key difference here is that we do not have the Lipschitz smoothness of the nonlinearity $F,$ which is important for the moving plane method used in [46]. Indeed, as we have already said, near $s=0,$ $F(s)\approx\sqrt{s}$; moreover, $F$ need not be smooth as $s\rightarrow 1$ (the maximum value of $\psi$) either. To deal with the singularity as $s\rightarrow 0,$ we use the fact that we have a lower bound on the growth of $\psi$ in the direction normal to the boundary, since the vorticity is nonzero (and constant) on the boundary. Then we invoke a refinement of the Hopf lemma for elliptic problems on small domains under some growth conditions on the coefficients. This effectively deals with the singularity as $s\rightarrow 0$. Remarkably, it turns out that the potential singularity as $s\rightarrow 1$ always has the right sign for the maximum principle (see the proof of Lemma 4.8).
1.5. Organization of the Paper
In Section 2, we establish some general results on analytic functions satisfying $\{f,g\}=0$ and recall a few results on zero sets of analytic functions. The main consequence of this is that, if an analytic Euler solution $\psi$ does not satisfy an equation of the type (1.4), we are ensured the existence of a level set of $\psi$ on which the gradient vanishes. In Section 3, we prove a few results about these sets, using more fundamentally that $\psi$ is an Euler solution. In particular, we show the analyticity of the non-isolated portions of the critical points with quadratic vanishing. This sets us up with an overdetermined elliptic problem in a subdomain of the original domain. At the end of Section 3, we then establish the rigidity of solutions in a domain whose boundary is a curve of critical points with vanishing greater than quadratic. In Section 4, we first establish the rigidity of solutions with a suitably chosen curve of critical points with quadratic vanishing, which is based on certain regularity properties of the function $F$ (defined in a suitable subdomain), and the moving plane method. Then, we finish the proof of Theorem 1.1 in Section 4.2 via a contradiction argument. In Section 4.3, we give a counterexample to Theorem 1.1 in an analytic non-simply connected bounded domain, hence showing the optimality of the statement.
2. Facts about $\{f,g\}=0$
@@
2.1. Key Facts
We start by giving a notation for the degree of vanishing of an analytic function at a point.
@@ 2.1. Key Facts
For any non-constant analytic function $f:\Omega\rightarrow\mathbb{R},$ and $x_{*}\in\Omega$, $D^{k}f(x_{*})$ denotes the tensor for the $k$-th order derivatives of $f$ at $x_{*}$, and we define $d_{f}(x_{*})$ to be the first $k\geq 1$ for which $D^{k}f(x_{*})\not=0$.
We then introduce the notations for the set of critical points of $f$ and those with even degree of vanishing:
@@
Assume that $\{f,g\}=0$ in some open connected set $\Omega\subset\mathbb{R}^{2}.$ Suppose that $\Gamma_{0}$ is an analytic curve connecting $x,y\in\Omega$ and that $\Gamma_{0}\cap\mathcal{C}_{e}(f)=\emptyset$. If $f(x)=f(y)$, then it must hold that $g(x)=g(y).$
Note that Lemma 1.3 is a direct corollary of Proposition 2.1. To prove this proposition, let us first establish a generalization of the well-known straightening lemma.
@@
The case $d=1$, which is that $\nabla f(x_{*})\not=0,$ is the standard straightening lemma, which is valid under the only assumptions that $f$ and $g$ are $C^{1}$ in a neighborhood of $x_{*}.$
@@
Strictly speaking, we do not need the generalization to $d>1$ for the results proven later; however, it is conceptually important in order to understand the difference between the “extendable” and “unextendable” domains later. The result may also be of independent interest. It should be noted that, in this generalization, analyticity plays a key role, with the statement not holding in the smooth class.
Proof of Lemma 2.2.
Without loss of generality, by applying suitable translations and a rotation, set $x_{*}=0$, $f(x_{*})=0,$ and $\partial_{1}^{d}f(0)>0.$ Assume that $d$ is odd. The point is that $\partial_{1}f(t,0)>0$ for $t\in[-\epsilon,0)\cup(0,\epsilon]$ for some $\epsilon>0$. In particular, the function $f(t,0)$ is real-analytic and invertible in $(-\epsilon,\epsilon)$. It follows that we can express its inverse using a Puiseux series for $t$ sufficiently small. In particular, there exists (a unique) $F$ that is analytic in the variable $s^{1/d}$ for which
$g(t,0)=F(f(t,0)),\,\,\text{for every}\,\,t\in(-\frac{\epsilon}{2},\frac{\epsilon}{2}).$
Let $\kappa$ be the radius of convergence of the Puiseux series of $F$ at $0.$ We now take $\delta<\frac{\epsilon}{2}$ be so small that $f(B_{\delta}(0))\subset(-\kappa,\kappa).$ We will show that
$g(x)=F(f(x)),\,\,\text{for every}\,\,x\in B_{\delta}(0).$
Fix any $x\in B_{\delta}(0).$ First, take $x_{1}\in(-\delta,\delta)\setminus\{0\}$. By assumption, $(x_{1},0)$ is not a critical point of $f$. It follows that there is a real analytic $G$ so that $g(z)=G(f(z)),$ for every $z$ in a neighborhood of $(x_{1},0).$ In particular, there is an $\epsilon(x_{1})>0$ such that for all $t\in(x_{1}-\epsilon(x_{1}),x_{1}+\epsilon(x_{1})),$ we have
@@
At this point, we have shown that there is a ball around each point of the segment $\{(t,0):t\in(-\epsilon,\epsilon)\setminus\{0\}\}$ where $g=F(f).$ Now we will extend this to the whole ball $B_{\delta}(0).$ Fix any $x\in B_{\delta}(0)$ with $x_{1}\neq 0$. Without loss of generality, assume $x_{2}>0.$ Now consider the function $g(x_{1},t),$ where $t\in[0,x_{2}].$ This is an analytic function of $t$, since $g$ is an analytic. Now, $F(f(x_{1},t))$ is analytic and equal to $g$ in a neighborhood of $t=0$. Moreover, $F(f(x_{1},t))$ is also analytic except at the possibly discrete set of points where $f(x_{1},t)=0$, where $F(f(x_{1},t))$ has a convergent Puiseux series (note that $f(x_{1},0)\not=0$). It follows that $F(f(x_{1},t))$ is actually analytic on $0,x_{2}]$ and that $g(x_{1},t)=F(f(x_{1},t))$ for every $t\in[0,x_{2}].$ In particular, $g(x_{1},x_{2})=F(f(x_{1},x_{2})).$
@@
For any point $\tilde{x}\in\Gamma_{0}$, by applying Lemma 2.2 to $f$ and $g$, there are a neighborhood $U_{\tilde{x}}$ and a Puiseux function $F_{\tilde{x}}$ such that $g=F(f)$ in $U_{\tilde{x}}$. As the collection of the open set $U_{\tilde{x}}$ forms an open cover for the compact set $\Gamma_{0}$, we can choose $U_{x^{1}},\cdots,U_{x^{m}}$, such that $U_{x^{i}}\cap U_{x^{i+1}}\neq\emptyset$ and $\Gamma_{0}\subset\cup_{i=1}^{m}U_{x^{i}}$. In the case where $f$ is constant in $\Gamma_{0}$, as the pair $(g,f)$ satisfies a functional relationship in each $U_{x^{i}}$, $g$ is constant in $U_{x^{i}}\cap\Gamma_{0}$. Notice that $\{U_{x^{i}}\}$ is a finite open cover for $\Gamma_{0}$, therefore, $g$ is constant in $\Gamma_{0}$.
In the case where $f$ is not constant in $\Gamma_{0}$, we claim that $F_{x^{1}}$ can be extended to be a Puiseux function $F$ in the range of $f(\Gamma_{0})$ with $g=F(f)$ in $\Gamma_{0}$. We first prove that $F_{x^{1}}$ can be extended to a Puiseux function $F$ in the range $f(U_{x^{1}}\cap\Gamma_{0})\cup f(U_{x^{2}}\cap\Gamma_{0})$, with $g=F(f)$ in $\Gamma_{0}\cap(U_{x^{1}}\cup U_{x^{2}})$. Here $F_{x^{1}}$ and $F_{x^{2}}$ are two Puiseux functions respectively in the range of $f(U_{x^{1}}\cap\Gamma_{0})$ and $f(U_{x^{2}}\cap\Gamma_{0})$, and the values of $F_{x^{1}}$ and $F_{x^{2}}$ agree in $f(\Gamma_{0}\cap U_{x^{1}}\cap U_{x^{2}})$. As $f$ is not constant in $\Gamma_{0}$, the analytic assumption of $f$ and $\Gamma_{0}$ implies that $f(\Gamma_{0}\cap U_{x^{1}}\cap U_{x^{2}})$ contains an open interval in $\mathbb{R}$. Now due to the fact that $f(\Gamma_{0}\cap U_{x^{1}})$ and $f(\Gamma_{0}\cap U_{x^{2}})$ are connected, the condition that $F_{x^{1}}$ and $F_{x^{2}}$ agree in $f(U_{x^{1}}\cap U_{x^{2}})$ implies that the values of $F_{x^{1}}$ and $F_{x^{2}}$ agree in $f(\Gamma_{0}\cap U_{x^{1}})\cap f(\Gamma_{0}\cap U_{x^{2}})$. Hence, we can extend $F_{x^{1}}$ with the aid of $F_{x^{2}}$ to a Puiseux function $F$ in $f(U_{x^{1}})\cup f(U_{x^{1}})$ and $g=F(f)$ in $\Gamma_{0}\cap(U_{x}^{1}\cup U_{x}^{2})$. Iterate the above procedure to extend $F_{x^{1}}$ to a Puiseux function $F$ in the range of $f(\Gamma_{0})$ such that $g=F(f)$ in $\Gamma_{0}$. Hence the proof of the proposition is completed. ∎
2.2. Zero Sets of Analytic Functions
In this section, we will briefly discuss the structure of the zero set of a real analytic function in two variables. Let us first recall a classical result concerning the structure of the zero set of a real analytic function.
@@ 2.2. Zero Sets of Analytic Functions
  3. (3)
    The nodal set $\{f=0\}$ near $x^{0}$ is a collection of an even number of analytic curves that branch out at $x^{0}$.
It is an immediate corollary of Proposition 2.3 that for any given point $x$, which is not isolated in the nodal set for a non-constant analytic function $f$, there are at least two analytic curves $\Gamma_{1}$, $\Gamma_{2}$ branching at $x$ and consisting of nodal points of $f$. In most cases, we cannot glue $\Gamma_{1}$ and $\Gamma_{2}$ together to form a single analytic curve. For example, in the case $f(x_{1},x_{2})=x_{1}^{2}-x_{2}^{3}$, the nodal set has a cusp at the origin. However, in the case where $\Gamma_{1}$ consists of critical points of the same degree (in particular, if the degree does not jump up at the common point), $\Gamma_{1}\cup\Gamma_{2}$ is an analytic curve.
@@
Let $x^{0}$ be a fixed point and we assume that there is an analytic curve $\Gamma_{1}:[0,1]\rightarrow\mathbb{R}^{2}$ with $\Gamma_{1}(0)=x^{0}$, and $d_{f}(x)=m\in[2,+\infty)$ for all $x\in\Gamma_{1}([0,1])$. Then the level set $\{x|f(x)=f(x^{0})\}$ near $x^{0}$ is an analytic curve $\Gamma_{0}$. Moreover, we have $\Gamma_{1}\subseteq\Gamma_{0}$ and for any point $x\in\Gamma_{0}$, $d_{f}(x)=m$.
@@
Without loss of generality, by applying a suitable rotation, we assume $x^{0}=0$, $f(0)=0$ and $\partial_{1}^{m}f(0)>0$. By applying the analytic inverse function theorem to $\partial_{1}^{m-1}f$ at 0, we have the set $\{\partial_{1}^{m-1}f=0\}$ near 0 is an analytic curve $\Gamma_{0}$ passing through 0. The existence of the analytic curve $\Gamma_{1}$ consisting of critical points of degree $m$ that starts at 0 ensures that for infinitely many points $\{x^{i}\}_{i=1}^{\infty}\subset\Gamma_{1}$ accumulate at $0\in\Gamma_{0}$, we have
(2.1) $D^{k}f(x^{i})=0,\,\,\text{for $1\leq k\leq m-1$ and $i\in\mathbb{N}$. }$
As $\Gamma_{0}$ and $f$ are analytic, it holds that
(2.2) $D^{k}f(x)=0,\,\,\text{for $x\in\Gamma_{0}$, and $1\leq k\leq m-1$.}$
Then due to the fundamental theorem of calculus, we have $f=0$ in $\Gamma_{0}$.
Now that we have identified an analytic curve $\Gamma_{0}$ passing through the origin, which is a level set of $f$ and contains $\Gamma_{1}$, it remains to find an open neighborhood $\mathcal{O}$ of the origin such that
$\{f=0\}\cap\mathcal{O}=\Gamma_{0}\cap\mathcal{O}.$
We notice the condition $\partial_{1}^{m}f(0)>0$ ensures that $(1,0)$ is not a tangent vector of $\Gamma_{0}$ at 0. Now, since $f$ and $\Gamma_{0}$ are analytic, we can choose $\mathcal{O}:=(-\epsilon,\epsilon)^{2}$ to be a small cube near 0 with the requirement that all points $x$ in $\mathcal{O}$ satisfy $\partial_{1}^{m}f>0$ and, in addition, for any point $x\in\mathcal{O}\setminus\Gamma_{0}$, there is a horizontal line $l_{x}$ connecting $x$ and a point $x_{\Gamma}$ in $\Gamma_{0}$. Now, as
$\partial_{1}f(x_{\Gamma})=\cdots=\partial_{1}^{m-1}f(x_{\Gamma})=0\quad\text{and}\quad\partial_{1}^{m}f>0\,\,\text{in}\,\,l_{x},$
applying the mean value theorem to $f$ in $l_{x}$, we have $f(x)\neq 0$. ∎
3. Structure of the Set of Critical Points